OpenStudy (anonymous):

What is the simplified form of 1/b - 1/a/1/b + 1/a

OpenStudy (anonymous):

Question:

OpenStudy (anonymous):

Answer choices:

OpenStudy (anonymous):

@Destinymasha @phi @Punkchick @Luigi0210 @Callisto @marylou004 @emily_grace2 Help me out please :)

OpenStudy (luigi0210):

Get common denominators first~

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

@tHe_FiZiCx99 @wolf1728 Can you help?

OpenStudy (luigi0210):

In order to combine fractions(when adding/subtracting) you have to get common denominators. so multiply each by a and b to get: \[\Huge \frac{(\frac{a}{a}*\frac{1}{b})-(\frac{b}{b}*\frac{1}{a})}{(\frac{a}{a}*\frac{1}{b})+(\frac{b}{b}*\frac{1}{a}})\]

OpenStudy (luigi0210):

Doing that without equation editor is frustrating .-.

OpenStudy (anonymous):

Thank you :) How would I get the answer?

OpenStudy (luigi0210):

Multiply it out and get: \[\Huge \frac{\frac{a}{ab}-\frac{b}{ab}}{\frac{a}{ab}+\frac{b}{ab}}\] So now you can add/subtract them: \[\Huge \frac{\frac{a-b}{ab}}{\frac{a+b}{ab}}\] Do you know what to do when you have a fraction divided by a fraction?

OpenStudy (anonymous):

Umm I kinda remember haha Multiply..something?

OpenStudy (luigi0210):

Uhm, yea. Take the bottom fraction, flip it and then multiply it: \[\Huge \frac{\frac{a-b}{ab}}{\frac{a+b}{ab}}=\frac{a-b}{\color{red}{ab}}*\frac{\color{red}{ab}}{a+b}\] And if you cancel the ab's \[\Huge\frac{a-b}{1}*\frac{1}{a+b}\] What do you get now?

OpenStudy (anonymous):

I don't know..do you flip it again and divide? I haven't done this stuff for months, I'm sorry

OpenStudy (luigi0210):

._. You just multiply >.> \[\huge \frac{a-b}{a+b}\]

OpenStudy (anonymous):

Oh okay haha That was too easy, I made it look difficult. Thank you so much!

OpenStudy (luigi0210):

You're welcome~